oblem 173 0 and falls with a A starts from alt of 4 m/N mrmtproblemID-5609325 Part A mass hari rallenen meter d summine the magnitude of the angular velocity at the answee with the appropriate ranits. Express your Palme. Units Part B of the normal component lasen ore meter the mag ace leration a point at the outer edge of the puhey Express your en with the appropriate
Answer #1

A)
given

wi = 0

a = 4 m/s^2

R = 100 mm = 0.1 m

angular acceleration of the pulley, alfa = a/r

= 4/0.1

= 40 rad/s^2

let t is the time taken for the hanging object to fall 1 m.

use, y = voy*t + (1/2)*a*t^2

y = 0 + (1/2)*a*t^2

==> t = sqrt(2*y/a)

= sqrt(2*1/4)

= 0.707 s

at t = 0.707 s,

w = wi + alfa*t

= 0 + 40*0.707

= 28.3 rad/s

B) Normal component of acceleration = R*w^2

= 0.1*28.3^2

= 80.1 m/s^2